# Microlens-enhanced SiPMs

### v1

@@ V-A Evaluation of the optical parameters
The MLAs have been evaluated at different production steps. Before the deposition of the lens on the silicon, a test replication on a glass wafer plate and a scanning electron microscope (SEM) image was produced to cross check the obtained dimensions. In Fig. 7 a lens height of $21.9\mu m$ is measured, where the nominal design parameter was $21.5\mu m$. For the distance between lenses the SEM image is produced from the top. Note that the characterisation of transparent non-flat optical components is difficult with an optical microscope as variable focal planes are required.
@@
The statistical measurements of the structures show an excellent correspondence for $R_{Lens}$ and $H_{Lens}$ as can be seen in Table IV. $R_{Lens}$ is measured indirectly via the measured minimal distance between two lenses (minimal gap) and the lens height is the height over the residual plane. The next steps of the implementation verification are the evaluation of the $RLT$ and the alignment of the replicated pattern with respect to the structures on the silicon. Unlike the optical parameter of the lens, these properties are not fixed for a given replication mold but depend on the replication process. It is therefore useful to measure for every chip the $RLT$ and validate the alignment at several points over the surface.
@@
The optical parameters extracted match very well the nominal values. As the simulation results of Table I show, a small difference for $H_{Lens}$ does not lead to important changes in efficiency. Misalignment is the most important factor that can decrease the efficiency of the MLA. Thus, a dedicated alignment and control system was implemented.
@@ V-B Alignment
The replication of the MLA is performed with a mold that requires a precise alignment with respect to the structures on the silicon. Markers were added to the silicon not only for the alignment process during replication but also for the quality assurance after MLA deposition. Microlenses with reduced diameter and height were integrated into the mold and placed within the alignment marker cross corners. The misalignment was measured with an optical microscope at every corner of the MLA via an automated process. The precision is typically better than $1\,\mu m$ in both the horizontal and vertical directions. A maximal misalignment of $2\,\mu m$ in diagonal was observed but for many samples the misalignment was of order of $1\,\mu m$ in diagonal. In Fig. 8 the relative position of the microlens implementation with respect to the alignment markers is illustrated.
Figure 8: (a) SEM image of two smaller size microlenses placed on top of the alignment marker. (b) Close view of the alignment markers filled with the microlenses. (c) Coated region with microlenses with coaxial illumination. (d) as for (c) but with diffuse illumination. The reflections of light on the top of the microlenses can be seen. The pixel structure is almost fully covered by the lenses and cannot be seen.
@@ VI Comparison with measurements
The evaluation of the MLA is based on a relative comparison between the MLA enhanced and the flat layer samples. The expected differences are a change in PDE and external optical cross-talk [^4] Photons exiting the silicon and reflected on the entrance window can produce external crosstalk. due to the different surface structures, although also the SPTR is expected to change. To extend the variation of samples, also bare silicon of the identical fabrication process was measured. The detectors are 128-channel arrays which yield data from many channels. This reduces the systematic uncertainties due to manufacturing differences and small defects of the microlens surface. The alignment is nevertheless linked to a 64-channel die and is expected to affect all channels on the same die.
@@ VI-A Breakdown voltage and gain measurement
The breakdown voltage is measured for every channel and even though they are integrated on the same silicon die, a spread is seen due to manufacturing differences. Over 128 channels it is typically in the range of $100-300\,mV$. The temperature dependence of $V_{BD}$ for this technology was measured to be $32\,mV/K$ at room temperature. An integrated Pt1000 temperature sensor on the detector module gives access to a precise temperature monitoring. With the temperature in the lab stabilized by air conditioning, variations were smaller than ±$\pm 1\,K$. To determine $V_{BD}$ and the gain (G), the bias voltage was increased in steps of $1\,V$ and the current $I_{Bias}$ recorded. At the same time, the pulse rates at thresholds of 0.5 and 1.5 photo-electron amplitude (PE) ($R_{0.5}$, $R_{1.5}$) were recorded with a high-speed oscilloscope. The gain at every $V_{BD}$ point is calculated as the ratio between the current and the sum of $R_{0.5\,PE}$ and $R_{1.5\,PE}$ and divided by the elementary charge (e) such that $G=I_{Bias}/(R_{0.5PE}+R_{1.5PE})/e$. A linear fit is used to extrapolate $V_{BD}$ and the slope is extracted from the fit to determine the gain. The results are shown in Fig. 9. The measurement is performed with light illuminating the detector to enhance the precision of the current measurement but with a sufficiently low pulse rate such that the random overlap of pulses in time is negligible. The measurement is repeated for every wavelength and the error is given as the error bars in the plot.
@@ VI-B Cross-talk and after-pulse
The ratio of the rates at thresholds $1.5PE$ and $0.5PE$ defines the cross-talk probability $P_{x-talk}=R_{1.5PE}/R_{0.5PE}$). The data were recorded from four detector channels each and the error bars indicate the spread between channels. $P_{x-talk}$ is increased by the flat and microlens coating. With an average thickness of $10\mu m$ for flat and $12.4\mu m$ for the microlens coating, and a pixel size of $41.7\mu m$, the external optical cross-talk is expected to increase over bare silicon. Even such a thin ($10-20\mu m$) coating increases the probability of external optical cross-talk. A reduction from $13\%@8V$ for flat to $8.5\%@8V$ for microlens, with almost identical $RLT$, is observed. This corresponds to a reduction in external optical cross-talk of $40\%@8V$. The cross-talk as a function of $\Delta V$ is given in Fig. 9.
At room temperature and $\Delta V<10\,V$, the after-pulse probability is below $0.5\%$. As the main results for PDE and light yield improvement are calculated as ratios, it is well justified to neglect after-pulse corrections.
@@ VI-C PDE with narrow, normally incident light source
The three different detectors were characterized with a narrow, normally incident light source. A Xenon lamp and a tunable monochromator are used to produce a monochromatic light beam that is coupled into a multimode glass fibre and homogenised with a microlens based homogeniser  [^5] https://www.lightsource.tech/en/accessories/homogenizer. The detector (SiPM array) together with a reference photodiode (PD) is mounted at a distance of $350\,mm$ on a motorised linear stage, which allows centering of the detectors with respect to the light beam. A homogeneity better than ±$\pm 0.5\%$ is achieved over the reference PD of $100\,mm^{2}$ whereas for the small size SiPM channel $0.408\,mm^{2}$ a much better value is reached. With the calibration curve for the PD (sensitivity as a function of wavelength) and the ratio of the surface between the two detectors, the absolute PDE is calculated with the recorded pulse rate $R_{0.5\,PE}$ corrected by the cross-talk rate $R_{1.5\,PE}$ and the dark pulse rate. The dark rates are typically $20-100$ times smaller than the rates with light. For the evaluation of the MLA, we provide absolute and relative change in PDE where the latter has a lower systematic uncertainty as it is not dependent on the absolute PD calibration nor the homogeneity of the illumination.
@@
The measurements are performed in a fully automated setup where a limited temperature control (by air conditioning) is used. The $V_{bd}$ is calculated for every measurement for every channel and the PDE is given as a function of a $\Delta V$ in Fig.10. The results show only small fluctuations between channels on the same silicon dies. The error bars indicated in the plots are RMS for four channels. The wavelength dependency can be explained by a change in the refractive index of the microlens polymer material which is between $1.53-1.51$ in the wavelength range between $400-500\,nm$. The PDE increase is lower for higher $\Delta V$ which is in accordance with the effect expected due to the low field region of the pixel. The simulated increase with normal incident light for this implementation is $27.91\%$ at low $\Delta V$ (with low field region) and $21.11\%$ at high $\Delta V$ (without low field region). This can be compared with the measurement result of $22\%$ at low $\Delta V$ and $17\%$ at high $\Delta V$.
@@ VI-D Comparison with SciFi light source
The microlens parameters were optimised for the SciFi as a light source and represent one of the goals of the MLA project. The test with the SciFi was performed at CERN SPS with pions at a momentum of $p=180\,GeV/c$. The SciFi module was arranged to perform a longitudinal scan. The distance between particle beam and photodetector was varied between $20\,cm-230\,cm$ and the light intensity from normal incident particles was measured. With a Gaussian fit, the median is extracted from the characteristic Landau distribution introduced by the energy deposit fluctuations. This value is called the light yield (LY) of the SciFi tracker. As a consequence of the mirror placed at the far end of the SciFi mat at a distance of $245\,cm$ from the SiPM, the light detected is composed of a direct component and a mirrored component. At a distance of $20\,cm$, the mirrored component is reduced to about $20\%$ due to attenuation. A fibre mat of $130\,mm$ width was equipped with four flat layer and four microlens dies, 256 channels each. In Fig. 11 the LY is plotted as a function of $\Delta V$ and given for two distances from the detector. The LY measurement has been performed after a gain calibration with a light injection system for each acquisition channel.
@@
For the SciFi light distribution for a propagation length of $230~{}cm$, the simulated LY enhancement is $24.57\%$ for low $\Delta V$ and $18.59\%$ for high $\Delta V$, and the measured values with the SciFi are $22.5\%$ for low $\Delta V$ and $20.5\%$ for high $\Delta V$. Note that the simulation was performed with a light exit angle distribution corresponding to a $230~{}cm$ long fibre.
@@ VII Conclusion
In this work, we present a novel implementation of MLA with the goal to enhance the PDE of SiPMs. The implementation was guided by the technology’s limitation of MLA deposit by a mold and a UV curable polymer. To optimise the most important parameters, the geometry of the SiPM pixel was simulated with three regions: the dead region, the low field region representing a reduced PDE region at low $\Delta V$ and the fully active region. The MLA consists of spherical lenses placed in a checkerboard pattern aligned with the pixel structure. The simulation enables the optimisation of the parameters for lens diameter, lens height, and residual layer thickness. MLAs at the wafer level were deposited on SiPM multichannel arrays with $41.7\,\mu m$ pixel size from FBK and compared with bare and flat layer coated devices. The performance was measured in the lab by injecting monochromatic narrow normal incident light and in a fibre tracker module with high energetic particles. The results are presented in detail as a function of wavelength and $\Delta V$ for the lab measurement and as a function of injection distance and $\Delta V$ for the SciFi measurement.
In summary, at $\Delta V=2\,V$, an increase of $22\%$ is observed for the narrow normal light source and $17\%$ at $\Delta V=8V$. For the SciFi light source, the increase is $22.5\%$ and $20.5\%$ at $\Delta V=2\,V$ and $\Delta V=6\,V$, respectively. These values correspond to a $10-15\%$ overestimation by the simulation. As the implementation depends on the precise alignment of MLA and the silicon structure as well as on the optical parameters of the implementation, the results of simulation and measurements can be considered in good agreement.
In addition to higher PDE, the MLA-enhanced detectors show also a reduction of $40\%$ of external cross-talk compared to a flat coated detector. An improvement of the timing performance in SPTR is expected due to the partial masking of the border region (deflection of the photons towards the pixel center by the lens) which was reported by other works and seen in preliminary measurements in our lab.
For low $\Delta V=2V$ the prospects of MLA enhanced SiPMs with the SciFi tracker technology are excellent. With the latest implementation of HPK $41.7\mu m$ pixel size, a simulated improvement of $39\%$ is expected. With the smaller $31.3\mu m$ pixel size from FBK $28\%$ improvement is predicted. These two configurations are currently in preparation for MLA deposition and new results will be available by the end of 2024.

### v2

@@ V-A Evaluation of the optical parameters
The MLAs have been evaluated at different production steps. Before the deposition of the lens on the silicon, a test replication on a glass wafer plate and a scanning electron microscope (SEM) image were produced to cross-check the obtained dimensions. In Fig. 7 a lens height of $21.9\mu m$ is measured, where the nominal design parameter was $21.5\mu m$. For the distance between lenses, the SEM image is produced from the top. Note that the characterisation of transparent non-flat optical components is difficult with an optical microscope as variable focal planes are required.
@@
The statistical measurements of the structures show an excellent correspondence for $R_{Lens}$ and $H_{Lens}$, as can be seen in Table IV. $R_{Lens}$ is measured indirectly through the measured minimum distance between two lenses (minimal gap), and the height of the lens is the height over the residual plane. The next steps of the verification of the implementation are the evaluation of the $RLT$ and the alignment of the replicated pattern with respect to the structures on the silicon. Unlike the optical parameter of the lens, these properties are not fixed for a given replication mould but depend on the replication process. It is therefore useful to measure for every chip the $RLT$ and validate the alignment at several points on the surface.
@@
The extracted optical parameters match the nominal values very well. As the simulation results in Table I show, a small difference for $H_{Lens}$ does not lead to important changes in efficiency. Misalignment is the most important factor that can decrease the efficiency of the MLA. Thus, a dedicated alignment and control system was implemented.
@@ V-B Alignment
The MLA replication is performed with a mold that requires a precise alignment with respect to the structures on the silicon. Markers were added to the silicon not only for the alignment process during replication but also for quality assurance after MLA deposition. Microlenses with reduced diameter and height were integrated into the mold and placed within the cross corners of the alignment marker. The misalignment was measured with an optical microscope in every corner of the MLA through an automated process. The precision is typically better than $1\,\mu m$ in both the horizontal and vertical directions. A maximum misalignment of $2\,\mu m$ in diagonal was observed, but for many samples the misalignment was of the order of $1\,\mu m$ in diagonal. In Fig. 8 the relative position of the microlens implementation with respect to the alignment markers is illustrated.
Figure 8: (a) SEM image of two smaller size microlenses placed on top of the alignment marker. (b) Close view of the alignment markers filled with the microlenses. (c) Coated region with microlenses with coaxial illumination. (d) With a segmented diffuse light source, the reflection on each microlens is visible and illustrates the curved surface. In this mode, the flat regions between lenses appear black and the pixel structure (routing lines) is made invisible due to the lenses.
@@ VI Comparison with measurements
The evaluation of the MLA is based on a relative comparison between MLA enhanced and flat layer samples. The expected differences are a change in PDE and external optical cross-talk [^1] Generated by photons from a primary avalanche exiting the silicon, reflected on the entrance window and producing a sub-sequent avalanche. due to the different surface structures, although also the SPTR is expected to change. To extend the variation of samples, we also measured bare silicon from the identical fabrication process. The detectors are 128-channel arrays which yield data from many channels. This reduces the systematic uncertainties caused by manufacturing differences and small defects in the microlens surface. However, the alignment is linked to a 64-channel die and is expected to affect all channels on the same die.
@@ VI-A Breakdown voltage and gain measurement
The breakdown voltage is measured for every channel and, although they are integrated on the same silicon die, a spread is seen due to manufacturing differences. Over 128 channels, it is typically in the range of $100-300\,mV$. The temperature dependence of $V_{BD}$ for this technology was measured to be $32\,mV/K$ at room temperature. An integrated Pt1000 temperature sensor on the detector module gives access to precise temperature monitoring. With the temperature in the lab stabilised by air conditioning, the variations were smaller than ±$\pm 1\,K$. To determine $V_{BD}$ and the gain (G), the bias voltage was increased in steps of $1\,V$ and the current $I_{Bias}$ recorded. At the same time, pulse rates at thresholds of 0.5 and 1.5 photo-electron amplitude (PE) ($R_{0.5}$, $R_{1.5}$) were recorded with a high-speed oscilloscope. The gain at every $V_{BD}$ point is calculated as the ratio between the current and the sum of $R_{0.5\,PE}$ and $R_{1.5\,PE}$ and is divided by the elementary charge (e) such that $G=I_{Bias}/(R_{0.5PE}+R_{1.5PE})/e$. A linear fit is used to extrapolate $V_{BD}$ and the slope is extracted from the fit to determine the gain. The results are shown in Fig. 9. The measurement is performed with light illuminating the detector to enhance the precision of the current measurement but with a sufficiently low pulse rate such that the random overlap of pulses in time is negligible. The measurement is repeated for every wavelength, and the error is given as the error bars in the plot.
@@ VI-B Cross-talk and after-pulse
The ratio of the rates at thresholds $1.5PE$ and $0.5PE$ defines the cross-talk probability $P_{x-talk}=R_{1.5PE}/R_{0.5PE}$). The data were recorded from four detector channels each, and the error bars indicate the spread between channels. $P_{x-talk}$ is increased by the flat and microlens coating. With an average thickness of $10\mu m$ for flat surfaces and $12.4\mu m$ for microlens coating, and a pixel size of $41.7\mu m$, the external optical cross-talk is expected to increase over bare silicon. Even such a thin ($10-20\mu m$) coating increases the probability of external optical cross-talk. A reduction from $13\%@8V$ for flat to $8.5\%@8V$ for microlens is observed, with almost identical $RLT$. This corresponds to a reduction in external optical cross-talk of $40\%@8V$. The cross-talk as a function of $\Delta V$ is given in Fig. 9. At room temperature and $\Delta V<10\,V$, the after-pulse probability is below $0.5\%$. As the main results for PDE and light yield improvement are calculated as ratios, it is well justified to neglect after-pulse corrections.
@@ VI-C PDE with narrow, normally incident light source
The three different detectors were characterised with a narrow, normally incident light source. A Xenon lamp and a tunable monochromator are used to produce a monochromatic light beam that is coupled into a multimode glass fibre and homogenised with a microlens based homogeniser [^2] https://www.lightsource.tech/en/accessories/homogenizer. The detector (SiPM array) together with a reference photodiode (PD) is mounted at a distance of $350\,mm$ on a motorised linear stage, which allows for centering of the detectors with respect to the light beam. A homogeneity better than ±$\pm 0.5\%$ is achieved over the reference PD of $100\,mm^{2}$ whereas for the small SiPM channel $0.408\,mm^{2}$ a much better value is reached. With the calibration curve for the PD (sensitivity as a function of wavelength) and the ratio of the surface between the two detectors, the absolute PDE is calculated with the recorded pulse rate $R_{0.5\,PE}$ corrected by the cross-talk rate $R_{1.5\,PE}$ and the dark pulse rate. The dark rates are typically $20-100$ times smaller than the rates with light. For the evaluation of the MLA, we provide absolute and relative changes in PDE, where the latter has a lower systematic uncertainty, as it is not dependent on the absolute PD calibration nor the homogeneity of the illumination.
@@
Measurements are performed in a fully automated setup where limited temperature control (by air conditioning) is used. The $V_{bd}$ is calculated for each measurement for every channel, and the PDE is given as a function of a $\Delta V$ in Fig.10. The results show only small fluctuations between channels on the same silicon dies. The error bars indicated in the plots are RMS for four channels. The wavelength dependence can be explained by a change in the refractive index of the microlens polymer material, which is between $1.53-1.51$ in the wavelength range between $400-500\,nm$. The increase in PDE is lower for higher $\Delta V$, which is in accordance with the effect expected due to the low field region of the pixel. The simulated increase with normal incident light for this implementation is $27.91\%$ at low $\Delta V$ (with a low field region) and $21.11\%$ at high $\Delta V$ (without a low field region). This can be compared with the measurement result of $22\%$ at low $\Delta V$ and $17\%$ at high $\Delta V$.
@@ VI-D Comparison with SciFi light source
The microlens parameters were optimised for the SciFi as a light source and represent one of the goals of the MLA project. The test with the SciFi was performed at CERN SPS with pions at a momentum of $p=180\,GeV/c$. The SciFi module was arranged to perform a longitudinal scan. The distance between the particle beam and the photodetector was varied between $20\,cm-230\,cm$ and the light intensity of the normal incident particles was measured. With a Gaussian fit, the median is extracted from the characteristic Landau distribution introduced by the energy deposit fluctuations. This value is called the light yield (LY) of the SciFi tracker. As a consequence of the mirror placed at the far end of the SciFi mat at a distance of $245\,cm$ from the SiPM, the detected light is composed of a direct component and a mirrored component. At a distance of $20\,cm$, the mirrored component is reduced to about $20\%$ due to attenuation. A fibre mat of $130\,mm$ width was equipped with four flat layer dies and four microlens dies, 256 channels each. In Fig. 11 the LY is plotted as a function of $\Delta V$ and given for two distances from the detector. The LY measurement has been performed after a gain calibration with a light injection system for each acquisition channel.
@@
For the SciFi light distribution for a propagation length of $230~{}cm$, the simulated LY enhancement is $24.57\%$ for low $\Delta V$ and $18.59\%$ for high $\Delta V$, and the measured values with SciFi are $22.5\%$ for low $\Delta V$ and $20.5\%$ for high $\Delta V$. Note that the simulation was performed with a light exit angle distribution corresponding to a $230~{}cm$ long fibre.
@@ VII Conclusion
In this work, we present a novel implementation of MLA with the objective of improving the PDE of SiPMs. The implementation was guided by the technology limitation of the MLA deposit by a mold and a UV-curable polymer. To optimise the most important parameters, the geometry of the SiPM pixel was simulated with three regions: the dead region, the low field region representing a reduced PDE region at low $\Delta V$ and the fully active region. The MLA consists of spherical lenses placed in a checkerboard pattern aligned with the pixel structure. The simulation enables the optimisation of the parameters for lens diameter, lens height, and residual layer thickness. MLAs at the wafer level were deposited on SiPM multichannel arrays with $41.7\,\mu m$ pixel size from FBK and compared with bare and flat layer coated devices. The performance was measured in the lab by injecting monochromatic narrow normal incident light and in a fibre tracker module with high energetic particles. The results are presented in detail as a function of wavelength and $\Delta V$ for the laboratory measurement and as a function of injection distance and $\Delta V$ for the SciFi measurement.
In summary, at $\Delta V=2\,V$, an increase of $22\%$ is observed for the narrow normal light source and $17\%$ at $\Delta V=8V$. For the SciFi light source, the increase is $22.5\%$ and $20.5\%$ at $\Delta V=2\,V$ and $\Delta V=6\,V$, respectively. These values correspond to an overestimation of $10-15\%$ by simulation. As the implementation depends on the precise alignment of the MLA and the silicon structure as well as on the optical parameters of the implementation, the results of simulation and measurements can be considered in good agreement.
In addition to the higher PDE, MLA-enhanced detectors also show a reduction of $40\%$ of external cross-talk compared to a flat coated detector. An improvement in timing performance in SPTR is expected due to the partial masking of the border region (deflection of the photons towards the pixel centre by the lens), which was reported in other works and seen in preliminary measurements in our lab.
The optical transparency of the less than $50\mu m$ thick MLA layer, in a wavelength range between $350nm$ and $700nm$, was measured before and after irradiation. No measurable change was observed with a neutron equivalent fluence of the expected radiation level in the LHCb upgrade 2 ($3.0\times 10^{12}~{}{\rm 1\,MeV\,n_{eq}/cm^{2}}$) and an ionising dose of $1kGy$. Further improvements by adding an anti-reflecting coating (ARC) to the polymer surfaces of the MLA can potentially decrease the Fresnel reflection loss. A prototype with optimised coating for the emission spectral range of the SciFi light source is foreseen. For low $\Delta V=2V$, the prospects of MLA enhanced SiPMs with SciFi tracker technology are excellent. With the latest implementation of HPK $41.7\mu m$ pixel size, a simulated improvement of $39\%$ is expected. The smaller $31.3\mu m$ pixel size from FBK $28\%$ is predicted to improve. The smallest pixel size for the replication method used is currently $16~{}\mu m$ and is limited by the precision of alignment and the thickness of the residual layer. These two configurations are currently in preparation for MLA deposition, and new results will be available by the end of 2024.